OpenStudy (anonymous):

A right triangle has a hypotenuse that is 3 feet longer than one leg. The other leg is 4 feet.

OpenStudy (anonymous):

did you draw a picture?

OpenStudy (anonymous):

I need to find out the two sides but I'm not sure where to put them into the pentagorus theoreum

OpenStudy (anonymous):

it doesn't have a picture

OpenStudy (anonymous):

pythagorean theorem... even if the problem doesn't have a picture you can always make one that uses the given information...

OpenStudy (anonymous):

ohhh i see sorry its because in brazil its called pentagoras lol and thanks now i see

OpenStudy (anonymous):

wait do i square both the x and the 3

OpenStudy (anonymous):

so using pythagoras' theorem, we get \[\left( x+3 \right)^2=x^2+4^2\] now you can solve for x...

OpenStudy (anonymous):

ok i got it thanks sooooo much!!!!

OpenStudy (anonymous):

you're welcome!